OpenStudy (anonymous):

(2x+5)(x-4) my calulater got 150, but I don't know how it worked the problem. Can someone show how to work this out?

OpenStudy (anonymous):

your calculator may not understand variables

OpenStudy (anonymous):

you should do four multiplications, sometimes denoted by the unfortunate mnemonic device "FOIL' for "first, outer inner last" or just multiply as you would for two 2 digit numbers

OpenStudy (anonymous):

I'll try that.

OpenStudy (anonymous):

good

OpenStudy (anonymous):

I haven't done foil in a long time. I think I messed up. so far i have 2x^2+3x/x=-1/3. can you show how to foil this?

OpenStudy (anonymous):

so far i have 2x^2-2x+5x+1= 2x^2+3x+1= 2x^2+3x=-1 2x^2+3x/3=-1/3 I don't know where to go from here

OpenStudy (anonymous):

Is your equation equal to zero?